26-30. The internal energy of a system increases by 20 J and the quantity of work done on a system is 50 J. Is the heat absorbed or given off? By how much?



1
Expert's answer
2021-03-29T05:16:33-0400

∆U = Q+ W

Increase = 20

W = -50

20 = q -50

Q = 20+50

= 70

The heat is absorbed